OpenStudy (anonymous):

How can I calculate w from following Equation: \(cot(2w) = \large \frac{w}{2}\)

OpenStudy (anonymous):

Please refresh question once..

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

just to check, you mean cot as in short for coth, the hyperbolic trig function?

OpenStudy (anonymous):

cot = cotangent right @waterineyes ?

OpenStudy (anonymous):

Yes, it is Cotangent..

OpenStudy (anonymous):

Ah, I see. We just have different terminology here in Australia. I've had a couple of embarrassing situations like this now. Afraid I can't help much then, sorry.

OpenStudy (anonymous):

I thought Mathematics is same everywhere.. :)

OpenStudy (anonymous):

mhm, same here?

OpenStudy (anonymous):

@ganeshie8 @.Sam. @AravindG

OpenStudy (anonymous):

Oh I meant I agree with you @waterineyes

OpenStudy (anonymous):

Most are the same, just that some terminology differs from country to country. It's similar to having two different text books, they can sometimes have different names, shortcuts, symbols for functions.

OpenStudy (anonymous):

Oh, cot and cotangent?? After all meaning is same buddy.. Here, we use \(cos^{-1}\) and other uses Arccos like that..

OpenStudy (anonymous):

So you want to solve for w, right?

OpenStudy (anonymous):

I tried, but i couldn't seem to get it :/

OpenStudy (anonymous):

Yes, @AravindG

OpenStudy (aravindg):

1 approach would be the graphical method

OpenStudy (aravindg):

I am thinking of the algebraic way.

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

I doubt that this might help...? \[\cot(2w) = \frac{ w }{ 2 } \rightarrow \frac{ 1 }{ \tan(2w) } = \frac{ w }{ 2 }\]

OpenStudy (anonymous):

@dan815 @Luigi0210

OpenStudy (anonymous):

In what sense, it can help??

OpenStudy (anonymous):

I do not know.

OpenStudy (aravindg):

I am not able to separate w completely to one side. What I got is this : \[\cot w-\tan w=2\cot2w\] \[\cot w-\tan w=2(\dfrac{w}{2})=w\]

OpenStudy (aravindg):

Wait a second ...Maybe you could form a quadratic in tan w or cot w from here and solve for it! I dont have time to work it here though..Best of luck!

OpenStudy (anonymous):

Suppose if I have a scientific calculator, then how can I put this equation in that calculator to find the value of \(\omega\)??

OpenStudy (anonymous):

@hartnn you are urgently required here..

OpenStudy (anonymous):

Newton's Method

OpenStudy (anonymous):

Hmmm, well \[ \cot (2w) = \frac{\cos(2w)}{\sin(2w)} \]So it does come out to be: \[ 2\cos(2w) = w\sin(2w) \] I don't see any definitive method here other than finding roots.

hartnn (hartnn):

which topic does this belong...i guess, no algebraic methods applicable

OpenStudy (abb0t):

have you tried in terms of cosine and sine?

OpenStudy (abb0t):

omfg. Lol. i Just saw someone suggested that. I'm such a meatball.

OpenStudy (anonymous):

I am not able to find it unless I use online calculators like wolfram alpha.. How will we eliminate w from the argument of cot??